# LAuReL: Learned Augmented Residual Layer

### v3

@@ 3.3 LAuReL-LR: Rank vs Accuracy
We note that for the LAuReL-LR version on the ResNet-50/ImageNet combination, there is a pattern in terms of the best accuracy achieved with different values of $r$. In the combined LAuReL-RW+LR version we experimented with different values of $r$, and computed the average of the best accuracy@1 achieved over 5 trials; see Figure 3. From Table 1, with the LAuReL-RW version alone we already achieve an average best accuracy@1 of 75.10%, therefore for the combined LAuReL-RW+LR version we would like to see the accuracy exceeding that.
We observe that when $r$ is small ($r\in\{4,8\}$), there is not a significant improvement over the baseline LAuReL-RW experiment. This could be because a very small $r$ acts as an information bottleneck in the low-rank product in (3). As $r$ increases, the accuracy reaches the maximum for $r=16,32$; beyond this, the accuracy seems to drop though still higher than the LAuReL-RW baseline.
We believe this unimodal phenomenon could be due to the number of parameters added to the model, which increases linearly in $r$, this would also require appropriate tuning of hyper-parameters such as the learning rate as well as the regularization penalty.
Another possible cause could be the initialization scheme for the low-rank matrices $A$ and $B$. In our experiments, we set $A$ to be initialized either using Xavier initialization for ResNet, or using an orthogonal initialization for the LLM experiments. $B$ was always initialized to zero. This is similar to the scheme used in Hu et al. (2022). We found that initialization made a significant difference in the performance of the LAuReL-LR variant, and we posit that further work studying and improving the initialization scheme of the low-rank matrices could lead to better performance of the variant.
Figure 3: Trend of accuracy as $r$ is varied in the LAuReL-RW+LR variant.
@@ 4 Efficiency
To achieve these favorable trade-offs, the above-mentioned models incorporate training techniques and architectural changes that can help improve quality while keeping training and inference cheap enough. Therefore, any proposed model efficiency technique including LAuReL should demonstrate not just improvement in quality, but also training and inference efficiency.
@@ 4.1.1 Number of Parameters
This is the most commonly quoted metric that is designed to capture the cost of training and serving the model. When the model in question is an MoE, there are two numbers quoted: the total number of parameters and the number of activated parameters. Note that a larger number of parameters implies greater forward and backward pass costs.
@@ 4.1.2 Latency (Training and Inference)
For training, a reasonable metric to consider would be the number of steps taken per second. Similarly, an important metric would be inference latency when deploying the model. For LLMs this can be broken down into the latency metrics in the warm-up stage (e.g., time-to-first-token, which can be further refined to prefill latency), and the generation stage (output tokens/second). For LLMs and other interactive models, both the time-to-first token and output tokens/second are important for a good user experience.
4.1.3 Memory
Peak memory used during training is another key metric that is tracked to ensure that accelerators can fit in the model graph and have enough free memory available for forward and backward passes. Inefficiencies can be compensated by strategies like rematerialization (Kumar et al., 2019), which can help reduce peak memory usage but at a cost of recomputing some activations. Similarly, peak memory used during inference is also an important metric to track.
@@ 4.2 Efficiency Analysis
Table 4: Analysis of extra params, memory, and latency incurred for each LAuReL variant, per instantiation, except the LAuReL-PA case where at any time no more than $k$ previous activations are in-memory, incurring a total $\Theta(kD)$ extra memory cost. Also note that the latency cost of LAuReL-LR can be tighter than $O(rD^{2})$, depending upon the exact matrix multiplication algorithm used. The latency complexities ignore the batch and sequence dimensions without loss of generality.
@@
(k denotes number of
of prev. activations)
LAuReL-LR+PA $2krD$ $\Theta(2krD$ $O(krD^{2}$
[…]
Table 4 lists the costs associated with using each LAuReL layer. We assume that we are working with a vector input of size $D$ (the last dimension of the input). This should not have any impact on the calculations or the generalization.
Note that the listed costs are per-LAuReL instantiation. Hence, if there are many transformer layers and if a particular LAuReL variant is used in each layer, then the relevant costs accumulate across the layers. We now examine the costs of the LAuReL variants.
LAuReL-RW is the cheapest with one or two scalars as parameters, a constant memory cost (the constant depending on the data type: fp32, fp16 / bf16, unit8 / int8, etc.), and a constant latency cost since the scalar multiplication might be optimized by the compiler.
The LAuReL-LR variant has $2rD$ parameters per instantiation ($rD$ for each of the two low-rank matrices, $A$ and $B$), and uses $\Theta(rD)$ memory. The latency is upper bounded by $O(rD^{2})$, since the input is a vector of dimension $D$ (ignoring the batch and sequence dimensions), and there are two matrix multiplications.
Finally, the LAuReL-PA variant uses $k$ previous activations. Assuming $h_{i}$ is an identity function, it uses $k$ extra parameters, and uses $\Theta(kD)$ extra memory for the previous activations. Since it requires $k$ extra vector additions, this introduces an additional latency of $O(kD)$.
@@ 4.3 Practical Recommendations
Given the experiments and the analysis of individual variants, LAuReL-RW would be the obvious first candidate to try. The LAuReL-RW+LR offers further improvement in model quality, and seems to provide the best trade-off in terms of model quality improvements and the overhead of parameters, latency, and memory. We validated this in ResNet and LLM experiments.
The LAuReL-RW+LR+PA variant also leads to improvements over LAuReL-RW+LR as the ResNet experiments show. The main caveat is the additional $\Theta(kD)$ memory, which might be undesirable. However, this can be mitigated to some extent by choosing a small value of $k$ or by changing the model sharding/rematerialization methods.
@@ 4.4 Comparison with Naive Model Scaling
One of the advantages of LAuReL to highlight is that it is competitive against naive scaling. Concretely, given a modest additional budget for parameters, latency, memory, allocating that budget to LAuReL (or its variants) is likely to produce more gains than naively using it in an existing model architecture.
@@
We observed this in the ResNet experiments (Section 3.1), where LAuReL variants match the performance of naive scaling while adding 2.6$\times$ fewer parameters, thus showing the Pareto-efficiency of LAuReL (vs naive scaling).
@@
However, in the LLM experiments in Section 3.2, the gains achieved using LAuReL variants are at the cost of $\sim$ 0.01–0.1% more parameters. A naive way to use this ’additional’ budget would be, say, increase the model dimension ($D$). However, this is may not be feasible due to hardware limitations and memory alignment issues. Another way to use this ’additional’ budget would be to increase the vocabulary size. However, the gains would be limited since this will include more tokens only from the tail of the distribution and may not contribute much to the model quality. Interestingly, Karpathy (2023) reported that the NanoGPT model (Karpathy, 2022) pre-training was sped up by $\sim$ 25% when the vocabulary size was made divisible by 64.
To summarize, not only did we demonstrate that LAuReL outperforms naive scaling, but also in transformer-style models, naively distributing the relatively small additional budget of LAuReL parameters might actually hurt performance and would require non-trivial systems engineering.
@@ 5 Related Work
Since our larger goal is to improve the training and inference efficiency of deep learning models, we briefly discuss some research directions aimed at improving model efficiency.
Architectural Changes: Our work is inspired by recent model architecture improvements such as LoRA (Hu et al., 2022) and AltUp (Baykal et al., 2023) amongst others. However, they are not directly relevant to LAuReL. Indeed, LoRA is designed to efficiently fine-tune large pre-trained models and it works directly on the model weight matrices level by introducing low-rank ‘adapter’ weights that are learned during the fine-tuning stage, while other model weights are held constant. In contrast, LAuReL works at the residual connection level, which likely spans multiple weight matrices involved in the function $f$; furthermore, it is applied during the pre-training stage.
AltUp (Baykal et al., 2023) is designed to replicate the quality improvements of a model with a large model dimension, without having to pay the additional cost. It operates at the transformer-block level, constructing parallel ‘lightweight’ transformer blocks to approximate the model dimension scaling effect. In contrast, LAuReL operates at the residual connection level and does not aim to replicate the dimension scaling effect.
Interestingly, LAuReL can be applied in conjunction with both LoRA (during fine-tuning) and AltUp (during pre-training and fine-tuning).
He & Hofmann (2024) proposes several changes to transformer blocks to help improve model convergence; however these proposals are limited to transformer blocks.
Compression Techniques: Model compression techniques (Buciluǎ et al., 2006) such as quantization (Krishnamoorthi, 2018; Jacob et al., 2018), including ternary networks (Li et al., 2016; Ma et al., 2024) are commonly used to reduce model size and inference latency. Similarly, pruning and model sparsity (Gale et al., 2019; Liu et al., 2019) techniques have also been explored and implemented in hardware.
Learning Techniques: Distillation (Hinton et al., 2014) is a popular technique for improving a smaller (student) model quality using “soft-labels” from a larger, impractical (teacher) model (Sanh et al., 2019). Some distillation variants propose learning intermediate representations as well (Zagoruyko & Komodakis, 2016; Kim et al., 2023). Other techniques include works like Stacking (Reddi et al., 2023) and RaPTr (Panigrahi et al., 2024), which progressively grow and train the network to achieve an improved model quality while reducing model training time.
@@ 6 Conclusion
In this paper we introduce the LAuReL framework, which is a novel architectural change and a generalization of the residual / skip connection aimed at improving the model quality without significantly increasing the model size or latency. We study three versions (LAuReL-RW, LAuReL-LR, LAuReL-PA) that can be mixed-and-matched together, as we show in our experiments.
Through experiments, we demonstrate the efficacy of replacing the conventional residual connection with LAuReL on both vision and language tasks, while also providing evidence for its advantages over naive model scaling methods. For future work, we would like to try to further improve LAuReL with new variants with better quality v/s model footprint trade-offs.

### v4

@@ 3.3 LAuReL-LR: Rank vs Accuracy
We note that for the LAuReL-LR version on the ResNet-50/ImageNet combination, there is a pattern in terms of the best accuracy achieved with different values of $r$. In the combined LAuReL-RW+LR version, we experimented with different values of $r$, and computed the average of the best accuracy@1 achieved over 5 trials; see Figure 3. From Table 1, with the LAuReL-RW version alone we already achieve an average best accuracy@1 of 75.10%, therefore for the combined LAuReL-RW+LR version we would like to see the accuracy exceeding that.
We observe that when $r$ is small ($r\in\{4,8\}$), there is not a significant improvement over the baseline LAuReL-RW experiment. This could be because a very small $r$ acts as an information bottleneck in the low-rank product in (3). As $r$ increases, the accuracy reaches the maximum for $r\in\{16,32\}$; beyond this, the accuracy seems to drop though still higher than the LAuReL-RW baseline.
We believe this unimodal phenomenon could be due to the number of parameters added to the model (which increases linearly in $r$), since this would also require appropriate tuning of hyperparameters such as the learning rate as well as the regularization penalty.
Another possible cause could be how the low-rank matrices $A$ and $B$ are initialized. For $A\in\mathbb{R}^{D\times r}$, we used the Xavier initialization for the ResNet experiment and a column orthogonal initialization for the LLM experiments [^2] We use $A_{i,j}=1/\sqrt{rD}$ if $i~{}\mathrm{mod}~{}r=j$ and 00 otherwise. while $B$ was always initialized to zero; this is similar to the scheme used in Hu et al. (2022). We found that initialization made a significant difference in the performance of the LAuReL-LR variant, and we posit that further work studying and improving the initialization scheme of the low-rank matrices could lead to better performance of the variant.
In terms of the tuning required, since $r\ll D$, and if $D=512,768,1024,\dots$ as in typical LLMs, this leaves a small range of discrete values for $r$ (unlike real-valued hyperparameters such as learning rate, weight decay, etc). In our experience $r\in\{32,48,64\}$ work well for LLMs.
Figure 3: Average best accuracy@1 vs the rank ($r$) in the LAuReL-RW+LR variant.
@@ 4 Efficiency
To achieve these favorable trade-offs, the above-mentioned models incorporate training techniques and architectural changes that can help improve quality while keeping training and inference costs low. Therefore, any proposed model efficiency technique including LAuReL should demonstrate not just improvement in quality, but also training and inference efficiency.
@@ 4.1.1 Number of Parameters
This is the most common metric designed to capture the cost of training and serving the model. A larger number of parameters implies larger forward and backward pass costs.
@@ 4.1.2 Latency (Training and Inference)
For training, a key metric to consider would be the number of steps taken per second. Similarly, another key metric would be the inference latency when deploying the model. For LLMs this can be broken down into the latency metrics in the warm-up stage (e.g., time-to-first-token, which can be further refined to prefill latency), and the generation stage (output tokens/second). For LLMs and other interactive models, both the time-to-first token and output tokens/second are important for a good user experience.
4.1.3 Peak Memory
Peak memory used during training is another key metric that is tracked to ensure that accelerators can accommodate the model graph and have enough free memory available for forward and backward passes. Inefficiencies can be compensated by strategies like rematerialization (Kumar et al., 2019), which can help reduce peak memory usage but need recomputing some activations. Similarly, inference-time peak memory usage is also a key metric to track.
@@ 4.2 Efficiency Analysis
Table 4: Analysis of extra parameters, memory, and latency incurred for each LAuReL variant, per instantiation, except the LAuReL-PA case where at any time no more than $k$ previous activations are in-memory, incurring a total $\Theta(kD)$ extra memory cost. Also note that the latency cost of LAuReL-LR can be tighter than $O(rD^{2})$, depending upon which matrix multiplication algorithm used. For simplicity, the latency bounds drop the batch and sequence dimensions.
@@
LAuReL-LR+PA $2rkD+k$ $\Theta(2rkD+k)$ $O(krD^{2}+kD)$
Table 4 lists the costs associated for each LAuReL layer. We assume that we are working with a vector of size $D$ (the last dimension of the input). Note that the listed costs are per-LAuReL instantiation. Hence, if there are many transformer layers and if a particular LAuReL variant is used in each layer, then the relevant costs accumulate across the layers. We now examine the costs of the variants.
LAuReL-RW is the cheapest, with one or two scalars as parameters, a constant memory cost (the constant depending on the data type: fp32, fp16 / bf16, unit8 / int8, etc.), and a constant latency cost since the scalar multiplication might be optimized by the compiler.
The LAuReL-LR variant has $2rD$ parameters per instantiation ($rD$ for each of the two rank-$r$ matrices, $A$ and $B$), and uses $\Theta(rD)$ memory. The latency is upper bounded by $O(rD^{2})$, since the input is a vector of dimension $D$ (ignoring the batch and sequence dimensions), and there are two matrix multiplication steps.
Finally, the LAuReL-PA variant uses $k$ previous activations. Assuming $h_{i}$ is the identity function, it uses $k$ extra parameters, and uses $\Theta(kD)$ extra memory for the previous activations. Since it requires $k$ extra vector additions, this introduces an additional latency of $O(kD)$.
4.3 Ablation Study of LAuReL Variants
To supplement the efficiency analysis in Section 4.2, we provide an ablation study of LAuReL variants on a small LLM pre-training baseline. We used the C4 corpus (Raffel et al., 2020) with $\sim$ 10B tokens, and a $4\times 4$ Google Cloud TPU v6e (Trillium) topology [^3] We expect similar results with a comparable GPU setup. for compute.
In order to simplify the comparison across many ablations (and also to avoid the noise in downstream evaluations at the 10B tokens scale), we report model performance using the test loss, a reasonable proxy for downstream model quality. For each model we report the number of parameters, test loss, peak memory reported by profiling tools, and the average step time. The last two metrics are proxies for the theoretical memory and latency bounds respectively, as mentioned in Section 4.2. Lower is better for all metrics.
We trained our main baseline with 24 layers (Baseline-24) and 157.2M parameters, along with a larger baseline with 28 layers (Baseline-28) and 179.2M parameters. We ran all the LAuReL variants and two combinations (RW+LR, RW+LR+PA) on top of Baseline-24. For the LAuReL-LR variants and its combinations, we picked $r=32$. Similarly for LAuReL-PA variant and its combinations, we chose $k=3$. Table 5 shows the results.
Table 5: Performance of LAuReL variants on a small LLM pretraining task. LAuReL experiments are based on top of Baseline-24. LAuReL variants achieve a better test loss than Baseline-28, while having fewer parameters, requiring lesser memory, and being faster in average step time.
Variant Params Test Peak Avg. Step
(M) Loss Memory Time
(GB) (sec)
Baseline-24 157.20 3.0159 11.65 0.095
Baseline-28 179.23 2.9963 13.23 0.105
LAuReL variants (24 layers)
LAuReL-RW 157.20 2.9557 11.93 0.095
LAuReL-LR 158.40 2.9624 12.29 0.098
LAuReL-PA 157.22 2.9512 12.55 0.100
LAuReL-RW+LR 158.40 2.9531 12.57 0.099
LAuReL-RW+LR+PA 160.83 2.9499 12.90 0.104
As seen in the results, all LAuReL variants have a lower test loss than the Baseline-24, while having a negligible impact on additional parameters, peak memory, or average step time. In fact, LAuReL variants perform better than even Baseline-28 in terms of the test loss while using much fewer parameters, lower peak memory, and lower average step time.
@@ 4.4 Practical Recommendations
Given the experiments and the analysis of individual variants and their combinations, LAuReL-RW is clearly the first candidate to try. LAuReL-RW+LR offers further improvements to model quality, and seems to provide the best trade-off in terms of quality improvements and the additional overhead. We validated this in the ResNet (Section 3.1) and LLM experiments (Sections 3.2.1, 3.2.2, 4.3).
The LAuReL-RW+LR+PA variant also leads to improvements over LAuReL-RW+LR. The main caveat is the additional $\Theta(kD)$ memory, which might need to be monitored. If it is a concern, it can be mitigated to some extent by choosing a small value of $k$ or by changing the model sharding/re-materialization schemes.
Given the above tradeoffs in loss, memory, step time, etc. we recommend trying the LAuReL variants in the following order: RW $\rightarrow$ LR $\rightarrow$ RW+LR / PA $\rightarrow$ RW+LR+PA.
@@ 4.5 Comparison with Naive Model Scaling
One of the advantages of LAuReL we would like to highlight is that it is competitive against naive scaling. Concretely, given a modest additional budget for parameters, latency, and memory, allocating that budget to LAuReL variants is likely to produce larger gains than using that budget in naive scaling methods such as additional layers. We first observed this in the ResNet experiments (Section 3.1), where LAuReL variants match the performance of naive scaling while adding 2.6$\times$ fewer parameters, thus showing the Pareto-efficiency of LAuReL (versus naive scaling).
@@
In the LLM experiments in Section 3.2, the gains achieved using LAuReL variants are at the cost of $\sim$ 0.01–0.1% more parameters. A naive way to use this ‘additional’ budget would be to increase the model dimension ($D$) such that we match the extra parameter budget. However, this is may not be feasible due to hardware limitations and memory alignment issues. Another way to use this ‘additional’ budget would be to increase the vocabulary size. However, the gains would be limited since this will include more tokens only from the tail of the distribution and may not contribute much to the model quality. Interestingly, Karpathy (2023) reported that for the NanoGPT model (Karpathy, 2022) pre-training was sped up by $\sim$ 25% when the vocabulary size was made divisible by 64. Therefore naive scaling might also hurt metrics like latency and memory by making the model setup suboptimal on the hardware.
Additionally, we also conducted a detailed study comparing LAuReL with naive scaling on the same LLM pre-training task as LLM-4B as mentioned in Section 3.2.2. The baseline model and the the LAuReL experiments are similar to their counterparts in Section 3.2.2, except a configuration change, which allows adding a layer. Both have 40 layers each, and are referred to as Baseline-40 and LAuReL-40 respectively.
The naive scaling baseline had 41 layers, and is referred to as Baseline-41 henceforth. In Table 6, we present the respective number of parameters and average step times of the three models when training (forward + backward pass). A lower average step time is better. For Baseline-41 and LAuReL, we report the delta in number of parameters and average step time when compared to the original baseline.
Table 6: Comparison between Baseline-40 (40 layers), Baseline-41 (41 layers), and LAuReL in terms of parameters and average step time (forward and backward pass).
Model Params Avg. Step Time
(B) (sec)
Baseline-40 4.400 1.65
Baseline-41 4.560 (+3.63%) 1.68 (+1.81%)
LAuReL-40 4.404 (+0.1%) 1.69 (+2.42%)
Note LAuReL-40 adds only +0.1% parameters and incurs a step time penalty of 2.42%. In terms of extra parameters added, LAuReL-40 adds 36$\sim$ 36$\times$ 36 × fewer parameters than the extra parameters added by Baseline-41. The latency of LAuReL-40 is slightly higher than Baseline-41 since each layer incurs minor additional overhead.
Table 7 shows the downstream quality of the baselines and LAuReL-40 on 10 tasks across Math, General Reasoning, Reading Comprehension, Translation, and Mutimodal domains. LAuReL-40 wins on all tasks, except WMT23 and DocVQA where it matches the baselines. It outperforms not just Baseline-40, but also Baseline-41, achieving between 2%–21% improvement over Baseline-40’s performance.
Table 7: LAuReL-40’s comparison with naive scaling. Results that provide statistically significant improvement over the Baseline-40 are highlighted in green and bold. The percentages in the last row of the LAuReL-40 section indicate the relative improvement over the baseline.
Math General Reasoning Reading Comprehension Translation Multimodal
Method Math MGSM MMLU Belebele BookQA WMT23 MMMU Coco-Cap DocVQA TextVQA
Baseline-40 14.20 20.29 48.83 57.92 47.11 67.72 33.77 97.29 66.87 60.86
(4.40B Params) $\pm$ 0.88 $\pm$ 3.16 $\pm$ 0.81 $\pm$ 3.42 $\pm$ 4.06 $\pm$ 0.20 $\pm$ 3.11 $\pm$ 4.41 $\pm$ 2.67 $\pm$ 2.86
Baseline-41 14.50 20.29 49.10 59.30 42.77 67.74 35.33 98.50 66.18 60.23
(4.56B Params) $\pm$ 0.9 $\pm$ 3.15 $\pm$ 0.82 $\pm$ 3.34 $\pm$ 4.15 $\pm$ 0.21 $\pm$ 3.12 $\pm$ 3.53 $\pm$ 2.76 $\pm$ 2.87
LAuReL-40 15.11 23.12 50.32 62.65 57.22 67.71 37.57 99.27 66.92 63.15
(4.404B Params) $\pm$ 1.01 $\pm$ 3.51 $\pm$ 0.82 $\pm$ 3.15 $\pm$ 3.81 $\pm$ 0.19 $\pm$ 3.10 $\pm$ 5.03 $\pm$ 2.65 $\pm$ 2.82
(+6.48%) (+13.94%) (+3.05%) (+8.16%) (+21.46%) (-0.02%) (+11.25%) (+2.03%) (+0.07%) (+3.76%)
We posit that naive scaling by adding an extra layer or two in a very deep network does not always lead to a corresponding improvement in performance out of the box. For such networks, we might require hyperparameter tuning (learning rate, weight decay) to counter overfitting, or training on more tokens to realize the expected theoretical model performance predicted by the scaling laws.
It is also well known that with deeper networks, the residual stream starts to play a crucial role in model convergence and quality. We suggest that this is the primary reason why LAuReL performs well. It augments the residual stream with learned components, allocating capacity for the network to learn the linear components of the input better.
@@ 5 Related Work
Residual Stream. DenseNet (Huang et al., 2017) connects every pair of layers in the network and hence in the basic variant of DenseNet, all the activations need to be in memory. This is prohibitively expensive for deep LLMs and other modern transformers. When introducing dense-blocks, all previous activations within the block need to be visible to any given layer within the block; this requires refactoring the model architecture into dense blocks.
On the other hand, LAuReL requires minimal changes. In fact, in LAuReL-PA, which is the most similar to DenseNet, we make three design choices to achieve memory efficiency and performance. First, each layer only looks at the $k$ past activations ($k=3$ seems sufficient in our experiments). Second, we use low-rank linear functions to further reduce memory usage due to activations. Third, we use learned scalars ($\gamma_{i},\gamma_{i-1},\dots$) to weigh the previous activations (which we found to be crucial in practice), whereas DenseNet assumes a simple sum of the previous activations.
He et al. (2016) introduce variants of residual connections with different types of ‘gating’, which look similar to the LAuReL-RW variant, except that they use a much larger number of parameters ($O(D^{2})$ per layer), where LAuReL-RW uses one or two extra parameters per layer. Highway Nets (Srivastava et al., 2015) is similar to LAuReL-RW but they also use $D^{2}+D$ parameters; furthermore, they incur additional latency due a full-rank matrix multiplication. Residual Gates (Savarese, 2016) also is similar to LAuReL-RW, except they use ReLU as the gating function. However, LAuReL is a more general formulation.
Architectural Changes. Our work is inspired by recent model architecture improvements such as LoRA (Hu et al., 2022) and AltUp (Baykal et al., 2023) amongst others. LoRA is designed to efficiently fine-tune large pre-trained models and it works directly on the model weight matrices level by introducing low-rank ‘adapter’ weights that are learned during the fine-tuning stage, while other model weights are held constant. In contrast, LAuReL works at the residual connection level, which likely spans multiple weight matrices involved in the function $f$; furthermore, it is applied during the pre-training stage.
AltUp (Baykal et al., 2023) is designed to replicate the quality improvements of a model with a large model dimension, without having to pay the additional cost. It operates at the transformer-block level, constructing parallel ‘lightweight’ transformer blocks to approximate the model dimension scaling effect. In contrast, LAuReL does not aim to replicate the dimension scaling effect.
Interestingly, LAuReL can be applied in conjunction with both LoRA (during fine-tuning) and AltUp (during pre-training and fine-tuning). LAuReL can also be enabled at the same time as parameter-sharing techniques.
@@ 6 Conclusion
In this paper we introduce the LAuReL framework, which is a novel architectural change and a generalization of the residual/skip connection aimed at improving the model quality without significantly increasing the model size or latency. We study three versions (LAuReL-RW, LAuReL-LR, LAuReL-PA) that can be mixed-and-matched together.
Through experiments, we demonstrate the efficacy of replacing the conventional residual connection with LAuReL on both vision and language tasks, while also providing evidence for its advantages over naive model scaling methods. For future work, we aim to further improve LAuReL by developing new variants with better trade-offs between quality and model footprint.
Acknowledgments
We thank Cenk Baykal, Erik Vee, Fotis Iliopoulos, Khoa Trinh, and Sushant Sachdeva (in alphabetical order) for many helpful discussions. We would also like to thank Andrew Tomkins for suggesting the name for the work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.